OpenStudy (ujjwal):

\[a(b-c)x^{2}+b(c-a)x+c(a-b)=0\] if the roots of given equation are equal, then a,b,c are in a)AP b)GP c)HP d)HP or GP

OpenStudy (experimentx):

\[ (b(c-a))^2 = 4a(b−c)c(a−b)\] Let's simplify this

OpenStudy (ujjwal):

I tried the same process.. Got some weird and really long equation.. Couldn't figure out the relation.

OpenStudy (anonymous):

i understood only that the two equal roots are 1 each!:))

OpenStudy (ujjwal):

how do they turn out to be 1?

OpenStudy (anonymous):

Just satisfy it. Observation :P

OpenStudy (anonymous):

just use 1 in x's place. the equation gets satisfied!!!!

OpenStudy (anonymous):

a,b,c are in HP.

OpenStudy (anonymous):

i can just give a clue: use the relations between the roots and coefficients!

OpenStudy (anonymous):

Exactly. :D

OpenStudy (ujjwal):

It has only only one answer and its GP.. @anusha.p you were helpful. The roots are 1 so, \[\frac{-b(c-a)}{2a(b-c)}=1\]proceeding further i get condition for HP..

OpenStudy (anonymous):

finally u got it:))))))

OpenStudy (ujjwal):

yeah, thanks @anusha.p , my normal approach was useless here..

OpenStudy (anonymous):

Hold on.. It simplifies to, 2ac = ab + bc. => ab , ac, bc are in AP. => 1/c , 1/b , 1/a are in AP ( Dividing each term by abc)

OpenStudy (anonymous):

Doesnt that mean a, b, c are in HP?

OpenStudy (ujjwal):

Yeah @siddhantsharan its in HP

OpenStudy (anonymous):

So why'd you say, "It has only only one answer and its GP"

OpenStudy (ujjwal):

I mistyped it.. Below it i have written, "proceeding further i get condition for HP"

OpenStudy (anonymous):

Ohhhhh. Sorry. Didnt see that. :P

OpenStudy (ujjwal):

They don't give me an option to edit answers.. And if i delete it and write it again, everything will seem too weird.. The 'flow of conversation' will be disturbed..

OpenStudy (anonymous):

HAha. True that. Its fineee ::)